((2-x)/(7x)) times (14/(x-2))=
Answer by stanbon(75887)

((2-x)/(7x)) times (14/(x-2))
----
Note: (2-x)/(x-2) = -1
----
Your Problem:
(2-x)/(x-2)*(14/7x)
= -1(2/x)
= -2/x
================
Cheers,
Stan H.